### v2

# TF-Restormer: Complex Spectral Prediction for Speech Restoration

Abstract
Speech restoration in real-world conditions is challenging due to compounded distortions such as clipping, band-pass filtering, digital artifacts, noise, and reverberation, and low sampling rates. Existing systems, including vocoder-based approaches, often sacrifice signal fidelity, while diffusion models remain impractical for streaming. Moreover, most assume a fixed target sampling rate, requiring external resampling that leads to redundant computations. We present TF-Restormer, an encoder-decoder architecture that concentrates analysis on input-bandwidth with a time-frequency dual-path encoder and reconstructs missing high-frequency bands through a light decoder with frequency extension queries. It enables efficient and universal restoration across arbitrary input-output rates without redundant resampling. To support adversarial training across diverse rates, we introduce a shared sampling-frequency-independent (SFI) STFT discriminator. TF-Restormer further supports streaming with a causal time module, and improves robustness under extreme degradations by injecting spectral inductive bias into the frequency module. Finally, we propose a scaled log-spectral loss that stabilizes optimization under severe conditions while emphasizing well-predicted spectral details. As a single model across sampling rates, TF-Restormer consistently outperforms prior systems, achieving balanced gains in signal fidelity and perceptual quality, while its streaming mode maintains competitive performance for real-time use. Anonymous code and demos are available at https://tf-restormer.github.io/demo.
1 Introduction
Speech enhancement (Ephraim & Malah, 1984; Pascual et al., 2017) has historically progressed through isolated sub-tasks with dedicated models such as denoising (Hu et al., 2020; Ho et al., 2020), dereverberation (Han et al., 2015; Wang & Wang, 2020), declipping (Mack & Habets, 2019), and bandwidth extension or super-resolution (Liu et al., 2022a; Lee & Han, 2021). In real-world settings, however, multiple distortions often coincide and are further compounded by digital distortions including lossy codecs (e.g., MP3, Ogg). These factors obscure magnitude and phase in the signal, making coherent, faithful speech restoration substantially more difficult.
This growing complexity has prompted a shift towards general speech restoration using generative models to handle diverse distortions (Liu et al., 2022b; Serrà et al., 2022). Vocoder (Kumar et al., 2019; Kong et al., 2020a)-based approaches reconstruct waveforms from compressed representations (Liu et al., 2022b; Andreev et al., 2022; Babaev et al., 2024). While such approaches improve perceptual quality, they discard phase information and treat speech as a semantic abstraction rather than a physical signal, producing output signals that deviate significantly from the input. In contrast, waveform-based generative models (Oord et al., 2016; Serrà et al., 2022) including GAN- and diffusion-based methods operate directly on the waveform, avoiding the abstraction bottleneck. Diffusion models (Welker et al., 2022; Richter et al., 2023) directly generate waveforms or spectra with improved fidelity, but their temporal compression or iterative sampling precludes streaming.
A further limitation of existing approaches lies in their assumption of a fixed target sampling rate. Restoration models, as well as dedicated super-resolution methods (Kim et al., 2024; Lu et al., 2025), invariably begin by resampling the input to the target rate before processing. This design choice simplifies model training but introduces critical drawbacks: every input must be converted to the highest rate regardless of its native bandwidth, leading to redundant computation from resampling. Moreover, supporting multiple output rates requires either training separate models or repeatedly upsampling and downsampling, both of which are inefficient and impractical in real-world applications. These limitations call for a universal framework that analyzes input bandwidth directly and flexibly synthesizes arbitrary output rates without external resampling.
On the other hand, for denoising and separation tasks, models have been developed primarily in the complex STFT domain (Choi et al., 2018; Hu et al., 2020). In particular, recent works introduced time-frequency (TF) dual-path models (Dang et al., 2022; Wang et al., 2023), alternating sequence modeling along time and frequency axes. This design has proved effective for enhancement and separation tasks by preserving the frequency structure, and it naturally supports sampling-frequency-independent (SFI) formulations (Paulus & Torcoli, 2022; Zhang et al., 2023), since frequency bins can be treated as sequences whose length scales with the input rate while maintaining a consistent STFT frame duration. However, despite this flexibility, existing models still assume matched input-output rates. Moreover, because TF dual-path models explicitly preserve fine spectral structure, their computational cost grows substantially with higher sampling rates, making it difficult to apply to super-resolution tasks directly.
Figure 1: Overview of TF-Restormer. The model adopts an encoder-decoder design: TF-Encoder performs input-bandwidth analysis through stacked time and frequency modules, while the TF-Decoder reconstructs missing high-frequency bands with lightweight processing and learnable extension queries.
These limitations suggest the need for a new design that retains the strengths of TF dual-path processing while overcoming the inefficiency of fixed input-output rates. To this end, we present TF-Restormer, an encoder-decoder architecture for robust speech restoration under diverse degradations. Inspired by masked autoencoders (MAE) (He et al., 2022), TF-Restormer concentrates heavy processing in a TF dual-path encoder that analyzes the input bandwidth, while a lightweight decoder reconstructs the missing high-frequency components through learnable extension queries (Figure 1) with cross-self attention mechanism (Gupta et al., 2023). This asymmetric design enables arbitrary input-output sampling rates without external resampling, which minimizes redundant computation.
In summary, our contributions are as follows:
  • We design an asymmetric encoder-decoder framework based on a TF dual-path Transformer with SFI-STFT: the encoder focuses on the input bandwidth, while the lightweight decoder extends high frequencies via learnable queries with a cross-self mechanism. This design enables general speech restoration, including super-resolution across arbitrary input-output rates without external resampling, with a shared SFI-STFT discriminator that supports unified adversarial training across diverse rates.
  • We improve robustness and practicality of TF dual-path processing by enhancing the frequency module with a projection-based spectral inductive bias and extending the time module to a causal variant for streaming, ensuring stable operation under extreme degradations as well as real-time applicability.
  • We propose the scaled log-spectral loss as an auxiliary spectral objective that replaces conventional $\ell_{1}$$\ell_{2}$ spectral objectives and complements perceptual (Babaev et al., 2024) and adversarial training (Mao et al., 2017). By selectively emphasizing reliably predictable regions, this loss stabilizes optimization under severe distortions while mitigating oversmoothing.
2 Related Work
Vocoder-based restoration
Vocoder systems typically project speech into Mel features (Liu et al., 2022b; Babaev et al., 2024) or learned representations (Koizumi et al., 2023; Li et al., 2024) and synthesize waveforms with neural vocoders (Oord et al., 2016; Kumar et al., 2019; Kong et al., 2020a). They often achieve “studio-like” perceptual quality but reduce fidelity, since intermediate features serve as perceptual cues rather than physical signals. Most also rely on temporally compressed U-Net structures (Pascual et al., 2017; Stoller et al., 2018), limiting real-time use. In contrast, we directly predict complex STFTs within a TF dual-path structure, retaining both magnitude and phase while enabling a streaming variant with minimal changes. Recent studies (Kaneko et al., 2022; Lu et al., 2025) further show that direct spectral prediction with adversarial training can rival vocoder-based generation.
Diffusion-based restoration
Another approach is diffusion-based generation from waveform (Kong et al., 2020b; Serrà et al., 2022; Welker et al., 2022; Scheibler et al., 2024) or STFT inputs (Lemercier et al., 2023; Richter et al., 2023). While effective in perceptual quality, inference remains costly: even fast samplers require multiple denoising steps, hindering real-time use. Moreover, while diffusion excels at sampling diverse modes of a distribution, according to  Babaev et al. (2024), speech enhancement typically seeks the most likely clean realization consistent with the observation, not a diverse set of alternatives. As a result, diffusion introduces unnecessary complexity for restoration or enhancement task. Therefore, we target a main-mode estimate via single-pass complex spectral prediction augmented with an adversarial loss for perceptual sharpness.
TF dual-path models
For denoising, dereverberation, and separation, TF dual-path models (Dang et al., 2022; Cao et al., 2022) alternate sequence modeling along time and frequency, preserving spectral structure and supporting SFI designs (Zhang et al., 2023). They show strong performance in enhancement (Cao et al., 2022; Lu et al., 2023; Chao et al., 2024) and separation (Wang et al., 2023; Saijo et al., 2024; Shin et al., 2025), where the speech signal remains largely preserved, so prior models used identical block designs for time and frequency without considering domain-specific differences. Under challenging degradations, we address this with a projection-based frequency module that injects spectral inductive bias for more robust high-frequency recovery. In addition, while prior dual-path models assumed matched input-output rates, we extend them with an encoder-decoder structure that supports super-resolution beyond fixed-rate restoration.
Audio super-resolution
Conventional bandwidth extension and super-resolution models (Liu et al., 2022a; Han & Lee, 2022; Kim et al., 2024; Lu et al., 2025) typically assume a fixed target rate, upsampling all inputs before processing to fill missing bands. While effective, this introduces redundant computation and ties each model to a single output rate, requiring repeated downsampling of outputs for different targets. In contrast, our encoder-decoder framework confines heavy processing to the input bandwidth and restores missing high-frequency bands through lightweight extension queries, improving efficiency. Unlike prior work restricted to predetermined rates, our design naturally supports user-specified outputs through the TF dual-path backbone, offering flexibility and enabling training with data from diverse native sampling rates.
3 TF-Restormer
3.1 SFI Input-Output Formulation
As an SFI model (Paulus & Torcoli, 2022), TF-Restormer addresses arbitrary input sampling rates $f_{E}$ by constructing the STFT with a constant frame duration (SFI-STFT). Unlike conventional SFI that assumes matched input-output rates, we introduce the first decoupled formulation, enabling inference at user-specified output rates $f_{D}$. Concretely, given an input $x\in\mathbb{R}^{1\times N_{E}}$ with sampling rate $f_{E}$, its STFT is $\mathbf{X}\in\mathbb{R}^{F_{E}\times T\times 2}$, where $F_{E}$ and $T$ are the number of frequency bins and frames. TF-Restormer then predicts $\mathbf{Y}\in\mathbb{R}^{F_{D}\times T\times 2}$ corresponding to an output $y\in\mathbb{R}^{1\times N_{D}}$ at sampling rate $f_{D}$, satisfying $f_{E}:f_{D}=(F_{E}-1):(F_{D}-1)$ under the assumption of consistent frame duration.
To ensure universal applicability across sampling rates, we adopt a 40 ms analysis window with a 20 ms hop. This choice is a common unit in speech analysis and, being integer multiples across typical rates, guarantees consistent STFT construction at $\{8,16,22.05,24,32,44.1,48\}$ kHz without requiring resampling. The maximum number of frequency bins $F_{\text{max}}$ is 961 for $f_{E}=48$kHz.
Figure 2: Overall architecture of TF-Restormer. With SFI-STFT and iSTFT, arbitrary input-output sample rates can be addressed in a single model. The encoder feature serves as both the input and the fixed key/value for the Freq. cross-self module in the TF-Decoder.
3.2 Analysis Encoder and Extension Decoder
TF-encoder for input analysis
As illustrated in Figure 1 and 2, the TF-Restormer is constructed with TF-encoder and TF-decoder. The TF-encoder is responsible for analyzing the speech component from the input signals $\mathbf{X}\in\mathbb{R}^{F_{E}\times T\times 2}$. Before the TF-encoder, the input complex representation $\mathbf{X}\in\mathbb{R}^{F_{E}\times T\times 2}$ is first projected to $C_{E}$ dimension by 2d convolution (Conv2D) layer with kernel size of (3,3), followed by layer normalization (LN) (Ba et al., 2016). Then, sinusoidal positional embeddings are added along the frequency axis (Freq. PE). In the encoder, the projected feature is alternately processed by freq and time modules $B_{E}$ times to capture speech component.
TF-decoder with extension query
Then, the encoder features $\mathbf{Z}\hskip-2.84526pt\in\hskip-2.84526pt\mathbb{R}^{F_{\hskip-0.85358ptE}\times T\times C_{\hskip-0.85358ptE}}$ are embedded to both the input of decoder by projection layer and key/value for cross-attention in the frequency cross-self module as shown in Figure 2. Then, the projected encoder features are padded with learnable extension query $\mathbf{q}_{\text{ext}}\hskip-2.84526pt\in\hskip-2.84526pt\mathbb{R}^{(F_{D}-F_{E})\hskip-0.85358pt\times\hskip-0.85358ptT\hskip-0.85358pt\times\hskip-0.85358ptC_{D}}$ whose values are shared across all the frames and processed by frequency and time module by $B_{D}$ times. In particular, the frequency module performs cross-attention based on encoder feature as key/value. Then, the complex STFT values $\mathbf{Y}\in\mathbb{R}^{F_{D}\hskip-0.85358pt\times\hskip-0.85358ptT\hskip-0.85358pt\times\hskip-0.85358pt2}$ are estimated from decoder features by Conv2D layer.
3.3 Asymmetric TF dual-path module
In TF dual-path modules, given the feature with shape of $\mathbb{R}^{T\hskip-0.56905pt\times\hskip-0.56905ptF\hskip-0.56905pt\times\hskip-0.56905ptC}$ where $F\hskip-2.84526pt\in\hskip-2.84526pt\{F_{E},F_{D}\}$, time modules process $F$ independent sequences with lengths of $T$ while frequency modules consider the feature as $T$ sequences with lengths of $F$ as illustrated in Figure 1. However, frequency bins are usually more static and structured sequence with a fixed length compared to time frames, exhibiting a relatively consistent structural roles. Therefore, we introduce a frequency module with projection-based inductive bias, explicitly encoding structural regularities across frequency bins for stable restoration.
Figure 3: Unit modules in TF-Encoder and TF-Decoder. The (a) time module is based on MHSA with RoPE while (b) the frequency encoder module is based on MHSA with frequency projection layer. (c) The frequency decoder module employs MHCA based on key/value from the encoder features
We adopt common structure for both time and frequency modules borrowing from TF-Locoformer (Saijo et al., 2024). As shown in Figure 3, it consists of two macaron-style (Lu* et al., 2019) convolution feed-forward network (ConvFFN) with Conv1D with kernel size of $K$ for capturing local contexts. In ConvFFN, the expansion factor is 3 with SwiGLU as hidden activation. Between ConvFFN modules, multi-head self-attention (MHSA) is used for global contexts with $H$ heads. The time module performs MHSA on temporal frames with rotary positional encoding (RoPE) (Su et al., 2024) to offer the relative positions while the frequency module applies MHSA on frequency bins with the frequency projection layer to induce the structural bias.
Frequency cross-self module
For frequency cross-self module, we replace the first F-ConvFFN in the frequency self module with multi-head cross-attention (MHCA) based on key-value from the encoder feature $\mathbf{Z}_{\text{enc}}$ while query is high-frequency padded region by $\mathbf{q}_{\text{ext}}$, inspired by cross-self attention (Gupta et al., 2023). Therefore, MHCA conditionally operates when $f_{E}<f_{D}$ and extension query is padded, otherwise, bypassed.
Attention with structural bias
Linformer (Wang et al., 2020) introduced linear projections of key-value to reduce the computations of attention, while MLP-Mixer (Tolstikhin et al., 2021) went further by replacing MHSA with static linear operations. Motivated by these insights, we incorporate a frequency linear projection (Fig. 3(c)) to impose an inductive bias for the structural consistency of frequency bins on top of the benefits of dynamic attention. Since frequency bins exhibit consistent characteristics, we share the same projection layer across all modules and key-value mappings. Formally, for each head $h=1,\dots,H$, given key-value $\mathbf{K}_{h,c},\mathbf{V}_{h,c}\in\mathbb{R}^{T\times F}$ at channel $c$, a learnable projection matrix $\mathbf{A}_{h}\hskip-2.27621pt\in\hskip-1.42262pt\mathbb{R}^{{F_{\text{max}}}\times F_{\text{proj}}}$ with dimension $F_{\text{proj}}$ and maximum bins $F_{\text{max}}$ is applied as
$\tilde{\mathbf{K}}_{h,c}$ $=$ $[\mathbf{K}_{h,c},\mathbf{O}]\mathbf{A}_{h}\in\mathbb{R}^{T\times{F}_{\text{proj}}},\quad 1\leq c\leq C_{h},$ (1)
$\tilde{\mathbf{V}}_{h,c}$ $=$ $[\mathbf{V}_{h,c},\mathbf{O}]\mathbf{A}_{h}\in\mathbb{R}^{T\times{F}_{\text{proj}}},\quad 1\leq c\leq C_{h},$ (2)
where $\mathbf{O}\in\mathbb{R}^{T\times(F_{\text{max}}-F)}$ is a zero-padding matrix.
Streaming mode with mamba
The modular design of TF dual-path model further enables a seamless extension to streaming mode by replacing the time module with Mamba (Gu & Dao, 2024) blocks. Refer to Appendix C for detailed model configurations.
4 Training
The model is trained by two phases of pretraining and adversarial training. The model is trained with $f_{D}$ randomly selected from {16, 24, 44.1, 48}kHz at each step by downsampling target speech signals from VCTK dataset (Yamagishi et al., 2019). Based on speech sources, we simulated noisy reverberant signals by convolving the room impulse response (RIR) and noise samples from the DNS dataset (Reddy et al., 2020). We then applied various digital distortions including codecs and downsampled the signal to the sampling rates $f_{E}$ of 8k or 16kHz, which are common in practical restoration condition (see Appendix B.1 for details).
4.1 Pretraining
Perceptual loss
Following Babaev et al. (2024), we incorporate a self-supervised learning (SSL)-based perceptual loss to stabilize adversarial training and encourage human-aligned quality. Specifically, extracting features from a pretrained SSL model for both the enhanced and clean waveforms, we minimize the mean-squared-error between these representations:
$\mathcal{L}_{\text{p}}(\theta)=\mathbb{E}_{m,n}\!\left[\,|\phi(g_{\theta}(x))_{m,n}-\phi(s)_{m,n}|^{2}\right],$ (3)
given that ${y}=g_{\theta}(x)$ is output of restoration model $g_{\theta}(\cdot)$ with parameters of $\theta$. $\phi(\cdot)_{m,n}$ denotes the $m$-th element of $n$-th frame from its feature map. We utilize WavLM-conv (Chen et al., 2022) as in the previous study (Babaev et al., 2024).
Proposed scaled log-spectral loss
Because the perceptual loss is restricted to 16 kHz and it is beneficial to guide spectral details to complement the looseness of perceptual loss, a previous work adopted an $\ell_{1}$ distance on the magnitude spectrum (Babaev et al., 2024). However, because TF-Restormer operates directly on the complex spectrum, the model can be explicitly supervised on both real and imaginary components in addition to the magnitude. Therefore, when denoting the STFT of target signal $s$ by $S_{m,tf}\hskip-1.42262pt=\hskip-1.42262pt|S_{r,tf}\hskip-1.42262pt+\hskip-1.42262ptjS_{i,tf}|$ and that of model’s predicted signal $g_{\theta}(x)$ by $Y_{m,tf}\hskip-1.42262pt=\hskip-1.42262pt|Y_{r,tf}\hskip-1.42262pt+\hskip-1.42262ptjY_{i,tf}|$, we can extend to the complex domain as $\mathcal{L}_{\ell_{1}}\hskip-1.42262pt(\theta)\hskip-1.42262pt=\hskip-1.42262pt\sum_{c\in\mathcal{C}}\alpha_{c}\hskip-2.27621pt\cdot\hskip-1.42262pt\mathbb{E}_{t,f}\hskip-0.85358pt\left[|Y_{c,tf}\hskip-1.42262pt-\hskip-1.42262ptS_{c,tf}|\right]$, where $\mathcal{C}=\{r,i,m\}$ denotes the component index set and $\alpha_{c}$ are component weights.
However, even with complex supervision, while some regions are relatively easy to predict and receive consistent gradients, severely degraded or missing high-frequency regions yield unstable gradients and drive the model toward oversmoothing, or averaging, effects (Babaev et al., 2024). To address this, we propose a scaled log-spectral loss selectively emphasizing well-predicted regions while preventing poorly predicted regions from dominating:
$\mathcal{L}_{{\rm s}}(\theta)=\sum_{c\in\mathcal{C}}\alpha_{c}\cdot\mathbb{E}_{t,f}\left[\,w_{tf}\,\log\!\left(1+\frac{|Y_{c,tf}-S_{c,tf}|}{w_{tf}}\right)\right],$ (4)
where $w_{tf}$ is scale factor that controls the relative scaling of gradient. The formulation $w\log(1+d/w)$ formulation ensures large gradients on smaller distance $d$ than $w$, preserving regions where phase and magnitude are already reconstructed well, while suppressing the influence of large deviations, avoiding the averaging common in $\ell_{1}$ or $\ell_{2}$ losses. For choosing the weight value $w_{tf}$, we observed that a distance $|Y_{c,tf}-S_{c,tf}|$ tends to be proportional to the scale of the corresponding source component $S_{m,tf}$. Therefore, $w_{tf}$ is empirically set to $E[S_{m,tf}]$ by averaging over the frames. We use $\alpha_{m}\hskip-1.42262pt=\hskip-1.42262pt0.6$, $\alpha_{r}\hskip-1.42262pt=\hskip-1.42262pt0.2$, and $\alpha_{i}\hskip-1.42262pt=\hskip-1.42262pt0.2$. Finally, when combined with the perceptual loss, our proposed objective becomes
$\mathcal{L}_{\rm pre}(\theta)=\lambda_{\rm p}\mathcal{L}_{\rm p}(\theta)+\lambda_{\rm s}\mathcal{L}_{\rm s}(\theta).$ (5)
where $\lambda_{*}$ denote loss weighting factors. Therefore, perceptual loss mainly focuses on largely deviated component while the scaled log-spectral more on well-predicted component, making them complementary to each other. For pretraining, we train the TF-Restormer with $\lambda_{\text{p}}\hskip-1.42262pt=\hskip-1.42262pt100$ and $\lambda_{\text{s}}\hskip-1.42262pt=\hskip-1.42262pt1$.
4.2 Adversarial Training
After pretraining the generator with $\mathcal{L}_{\text{pre}}$, we introduce an adversarial loss component to reduce the artifacts and predict severely distorted or missing components. For adversarial training, we attach multi-scale STFT discriminators (Défossez et al., 2023) as $i$-th discriminator of $\varphi_{i}$ and apply least square GAN (LS-GAN) loss (Mao et al., 2017). For generator, generator LS-GAN and feature-matching loss (Kumar et al., 2019) terms are added, respectively:
$\mathcal{L}_{\rm gen}(\theta)\hskip 7.11317pt$ $=$ $\lambda_{\rm g}\mathcal{L}_{\rm g}(\theta)+\lambda_{\rm fm}\mathcal{L}_{\rm fm}(\theta)+\lambda_{\rm p}\mathcal{L}_{\rm p}(\theta)+\lambda_{\rm s}\mathcal{L}_{\rm s}(\theta)+\lambda_{\rm hf}\mathcal{L}_{\rm hf}(\theta),$ (6)
$\mathcal{L}_{\rm disc}(\varphi_{i})$ $=$ $\mathcal{L}_{\rm d}(\varphi_{i}),\quad i=1,...,I.$ (7)
where $\mathcal{L}_{\text{hf}}=\mathcal{L}_{\text{pesq}}+10\cdot\mathcal{L}_{\text{utmos}}$ is additional human-feedback loss (Babaev et al., 2024) for aesthetic quality with differentiable PESQ loss [^1] https://github.com/audiolabs/torch-pesq and UTMOS loss (Saeki et al., 2022). We performed adversarial training using $\mathcal{L}_{\text{gen}}(\theta)$ with $\lambda_{\text{g}}\hskip-1.42262pt=\hskip-1.42262pt0.005$, $\lambda_{\text{fm}}\hskip-1.42262pt=\hskip-1.42262pt0.1,\lambda_{\text{p}}\hskip-1.42262pt=\hskip-1.42262pt100$, $\lambda_{\text{s}}\hskip-1.42262pt=\hskip-1.42262pt1$, and $\lambda_{\text{hf}}\hskip-1.42262pt=\hskip-1.42262pt0.0001$. Notably, we assign small weights to $\mathcal{L}_{\text{g}}$ and $\mathcal{L}_{\text{hf}}$ to avoid excessive generation artifacts. (See Appendix B.2 for details of training configuration.)
Proposed multi-scale SFI-STFT discriminators
In conventional adversarial training, a dedicated generator for each target sampling rate is trained with a corresponding discriminator as well (Défossez et al., 2023; Babaev et al., 2024; Ju et al., 2024). For stable training of a single generator across diverse rates, we propose a discriminator based on SFI-STFT, which preserves a consistent physical frame duration across sampling rates. Implemented with strided Conv2D layers, STFT discriminator (Défossez et al., 2023) produces two-dimensional maps that provide local real/fake supervision in the time-frequency plane. This design maintains sensitivity to spectral structure while remaining agnostic to absolute frequency resolution, thereby supporting adversarial training across different rates without redundant resampling or multiple discriminators. We employ 5 SFI discriminators with STFT window durations of {20, 40, 60, 80, 100} ms.
5 Evaluation
5.1 Dataset and Metrics
UNIVERSE data for speech restoration
As speech restoration model, we evaluate on 100 synthetic samples generated by UNIVERSE authors (Serrà et al., 2022) to ensure comparability to prior works in $f_{E}=f_{D}=16$kHz setting. The dataset introduces diverse simulated degradations such as bandpass filtering, reverberation, codec compression, and transmission artifacts.
VCTK-DEMAND for denoising
We additionally evaluated the well-known Valentini denoising dataset (Valentini-Botinhao et al., 2017) for direct comparison with conventional enhancement models as speech enhancement benchmarking. The evaluation set (824 utterances) consists of noisy mixtures from two speakers under four SNR conditions (17.5, 12.5, 7.5, and 2.5 dB).
VCTK for super-resolution
For super-resolution evaluation, we construct paired data by downsampling 48 kHz clean utterances from the VCTK-0.92 dataset (Yamagishi et al., 2019). Beyond the clean case, we also create noisy-distorted conditions by adding degradations such as noise, reverberation, band-pass filtering, and codec effects, enabling a comprehensive evaluation of speech restoration with super-resolution. Note that the training simulation follows a similar procedure, which may provide a slight advantage to our model. (See Appendix D for details of the simulation.)
For the evaluation, we adopt non-intrusive perceptual estimators for mean opinion score (MOS): DNSMOS (Reddy et al., 2022), UTMOS (Saeki et al., 2022), and WVMOS (Andreev et al., 2022) to assess the perceptual quality regardless of reference signals. We also employ perceptual evaluation of speech quality (PESQ) (Rix et al., 2001) to assess the perceptual signal fidelity of restored signal compared to the reference. As a complement of PESQ based on lower band, we consider signal-to-distortion ratio (SDR) (Roux et al., 2018), log-spectral distance (LSD), and mel-cepstral distortion (MCD) (Fukada et al., 1992) to assess the full-band signal (44.1/48kHz). In addition, to evaluate the reference-aware speech generation quality by capturing semantic congruence, we report SpeechBERTScore(sBERT) and SpeechTokenDistance(sTokDis) (Saeki et al., 2024). We also confirm effectiveness on real recordings based on DNSMOS, UTMOS, and WVMOS (Appendix F).
5.2 Comparison with existing models
For the UNIVERSE dataset, we consider VoiceFixer (Liu et al., 2022b) as a Mel vocoder-based baseline, StoRM (Lemercier et al., 2023), UNIVERSE (Serrà et al., 2022), and UNIVERSE++ (Scheibler et al., 2024) as diffusion-based baselines, TF-Locoformer as a recent TF dual-path Transformer model, and FINALLY (Babaev et al., 2024) as a latest strong Mel-vocoder method. As shown in Table 1, VoiceFixer improves MOS but sacrifices fidelity due to its Mel representation, while FINALLY achieves the highest perceptual quality yet lacks signal fidelity, a trend confirmed in Table 2. Diffusion-based methods yield more balanced results by directly operating in the waveform or complex STFT. TF-Locoformer preserves signal-level fidelity but suffers from residual perceptual artifacts and failure to recover lost details and naturalness (MOS, LSD). In contrast, TF-Restormer provides consistent improvements under fidelity, semantics, and perceptual quality, with its streaming variant maintaining competitive effectiveness under causal constraints. This indicates its robustness across diverse degradations in a universal restoration setting. Note that all the compared models are offline methods.
[…]
Model Signal fidelity Semantic fidelity Non-intrusive quality
[…]
Input 1.55 5.58 1.89 10.21 0.84 0.69 1.76 2.19 2.23
Ground Truth 4.50 $\infty$ 0.00 0.00 1.00 1.00 4.28 4.26 3.33
VoiceFixer 1.77 -5.68 1.49 10.50 0.84 0.71 3.28 2.83 2.99
StoRM 1.76 9.01 1.67 6.87 0.84 0.70 3.14 2.70 2.94
[…]
TF-Locoformer 2.13 11.61 2.00 6.26 0.89 0.76 3.20 2.95 2.86
[…]
TF-Restormer 2.30 11.12 1.45 5.08 0.91 0.80 4.34 4.08 3.25
TF-Restormer-streaming 2.00 8.89 1.47 6.01 0.87 0.74 3.93 3.77 3.14
Next, we evaluate TF-Restormer on the VCTK+DEMAND focusing on denoising. In Table 2, we compare against DB-AIAT (Yu et al., 2021), MP-SENet (Lu et al., 2023), and TF-Locoformer as dedicated denoising models, and VoiceFixer, UNIVERSE, and FINALLY as universal restoration baselines. Since the input speech is already well preserved and only corrupted by additive noise, it favors models that minimize unnecessary generation and faithfully retain the input signal. Accordingly, dedicated denoising models outperform universal restoration models in terms of signal fidelity, as they are optimized to suppress noise without altering intact regions. In contrast, restoration models risk degrading reliability by over-modifying clean inputs, making them less trustworthy for such simple cases. While not surpassing dedicated denoising models in raw signal metrics, TF-Restormer achieves more consistent semantic and perceptual gains, showing strong generalization despite being designed for universal restoration.
Table 2: Results on VCTK-DEMAND for denoising task.
Model Signal fidelity Semantic fidelity Non-intrusive quality
[…]
Input 1.98  8.56 1.27 5.40 0.91 0.82 3.01 2.90 2.45
Ground Truth 4.50 $\infty$ 0.00 0.00 1.00 1.00 4.52 4.07 3.16
DB-AIAT 3.27 21.30 0.90 1.77 0.95 0.87 4.39 3.83 3.13
MP-SENet 3.61 21.03 0.85 1.58 0.95 0.88 4.35 3.86 3.12
TF-Locoformer 3.30 23.82 0.92 3.58 0.95 0.87 4.66 3.93 3.20
VoiceFixer 2.40 -1.12 0.97 7.40 0.90 0.81 4.15 3.50 3.08
UNIVERSE 2.84 18.77 1.17 2.20 0.92 0.83 4.32 3.75 3.03
FINALLY 2.94   4.60 - - - - 4.87 4.32 3.22
TF-Restormer 3.41 19.45 0.75 1.54 0.95 0.88 4.75 4.14 3.14
TF-Restormer-streaming 2.89 16.43 0.85 2.16 0.93 0.84 4.56 4.05 3.09
Finally, we experiment on the super-resolution task in Table 3, using a single model that directly supports arbitrary output sampling rates. For clean cases, we compare against dedicated super-resolution models: NVSR (Liu et al., 2022a), Frepainter (Kim et al., 2024), and AP-BWE (Lu et al., 2025), as well as VoiceFixer as a universal restoration baseline. As in Table 2, since the low-band of the input speech remains intact, dedicated models that concentrate on reconstructing the upper bands are favored. Unlike conventional approaches that rely on fixed input-output rates and often require zero-padding or redundant resampling, TF-Restormer leverages extension queries to dynamically expand the spectrum. With this versatility, TF-Restormer shows stable performance comparable to the dedicated models, faithfully retaining clean low-frequency regions while effectively generating high-frequency components. In adittion, under noisy-distorted conditions, TF-Restormer simultaneously restores corrupted regions and reconstructs missing high bands, demonstrating robust generalization beyond pure super-resolution. Overall, these results suggest the advantage of our model as a universal restoration framework that achieves bandwidth extension without sacrificing signal fidelity or requiring explicit resampling.
[…]
Method 8kHz $\rightarrow$ 16kHz 8kHz $\rightarrow$ 24kHz 8kHz $\rightarrow$ 44.1kHz 16kHz $\rightarrow$ 48kHz
[…]
clean (super-resolution only)
Input 2.53 1.84 3.33 2.91 2.03 3.33 3.44 2.44 3.33 3.17 1.31 3.83
[…]
TF-Restormer 0.89 1.29 4.10 0.95 1.48 4.09 1.01 1.74 4.10 0.97 1.28 4.11
Noisy-distorted (general speech restoration + super-resolution)
Input 3.36 11.38 1.73 3.49 11.60 1.73 3.64 11.47 1.73 3.48 11.37 1.73
[…]
StoRM 1.76 4.57 2.57 - - - - - - - - -
UNIVERSE++ 1.79 5.28 2.49 - - - - - - - - -
TF-Restormer 1.16 2.78 4.22 1.21 2.97 4.22 1.18 3.08 4.21 1.18 2.86 4.22
TF-Restormer-streaming 1.30 3.93 3.96 1.31 4.01 3.99 1.30 4.05 4.02 1.26 3.86 4.05
Table 4: Ablation Study. VCTK-C and -ND denote clean and noisy-distorted input from VCTK data in Table 3.
[…]
UNIVERSE (general restoration)
Mag.-only 2.07 6.03 3.82
Mag.+Cplx. 2.23 5.70 3.76
Scaled log 2.29 4.96 4.10
VCTK+DEMAND (denoising)
Mag.-only 2.93 3.13 3.95
Mag.+Cplx. 2.97 2.98 3.87
Scaled log 3.41 1.54 4.14
VCTK-C (super-resol., 8 $\to$ 16kHz)
Mag.-only 3.42 2.10 4.10
Mag.+Cplx. 3.48 1.86 4.07
Scaled log 3.70 1.29 4.10
(a) Effect of scaled log-spectral loss
[…]
VCTK-C (super-resol., 8 $\to$ 16kHz)
Dec-only. 2.12 2.93 3.82
w/o MHCA 1.04 1.81 4.07
w/ MHCA 0.89 1.29 4.10
VCTK-C (super-resol., 8 $\to$ 44.1kHz)
Dec-only. 3.25 10.48 3.36
w/o MHCA 1.35 1.70 4.07
w/ MHCA 1.01 1.74 4.10
VCTK-ND (super-resol., 8 $\to$ 16kHz)
Dec-only. 2.23 4.39 3.77
w/o MHCA 1.20 3.30 4.11
w/ MHCA 1.16 2.78 4.22
(b) Decoder design
[…]
UNIVERSE (general restoration)
w/o F-proj. 2.26 5.56 3.90
w/ F-proj. 2.29 4.96 4.10
VCTK+DEMAND (denoising)
w/o F-proj. 3.21 0.85 4.03
w/ F-proj. 3.41 0.75 4.14
VCTK-C (super-resol., 8 $\to$ 16kHz)
w/o F-proj. 3.54 1.68 3.97
w/ F-proj. 3.70 1.29 4.11
(c) Effects of frequency projection
5.3 Ablation study
To validate the effects of the proposed methods, we conduct an ablation study on scaled log-spectral loss, decoder design, and frequency projection module.
Effects of scaled log-spectral loss.
First, we compare the scaled log-spectral (S-log) loss against two other regression losses as a auxiliary objective to the perceptual loss: $\ell_{1}$ loss on magnitude of STFT (Mag.-only) as in (Babaev et al., 2024) and $\ell_{1}$ loss on real, imaginary, and magnitude (Cplx.) for STFT. In Table 4(a), using a complex loss shows the better results than only guiding magnitude component for all the cases. Also, using the proposed scaled log-spectral loss further improved the performance by stabilizing the gradients, confirming that finer spectral detail can be preserved without sacrificing global fidelity.
Decoder design.
Next, we validate our encoder-decoder design when $f_{E}\neq f_{D}$. Compared to the proposed method with MHCA, we test two variants: (i) a large decoder without an encoder, where extension queries are padded at the input, and (ii) an encoder-decoder without MHCA, where the decoder shares the same block design as the encoder. All three cases converge to similar behavior when $f_{E}=f_{D}$. For fairness, we rescaled the decoder-only case to match computation and memory. Illustrations are provided in Appendix E. As shown in Table 4(b), the encoder-decoder design yields more stable improvements by focusing computation on the input bandwidth. The decoder-only case produces unstable results in full-band reconstruction ($8\to 44.1$kHz). Adding the cross-self frequency module further improves performance: high-frequency bands are reconstructed more reliably via cross-attention using encoder features, whereas the isotropic encoder-decoder relies only on self-attention and thus performs worse.
Effects of frequency projection.
Finally, we examine the effect of the frequency projection module on robustness. Without projection, RoPE provides only positional context, resulting in the conventional symmetric TF dual-path structure. In contrast, adding projection to F-MHSA imposes structural bias along the frequency axis, yielding more stable training and consistent improvements across tasks, particularly under severe distortions and during high-frequency extension. These results demonstrate that structural bias is valuable for robust restoration under extreme conditions.
6 Conclusion
We presented TF-Restormer, a speech restoration model with an encoder-decoder design that enables efficient operation across arbitrary input and output sampling rates. By concentrating on the input bandwidth with a strong TF dual-path encoder and extending high frequencies through lightweight decoder queries with a cross-self mechanism, TF-Restormer achieves balanced improvements in both signal- and semantic-level fidelity while also showing robust performance on single-task benchmarks such as denoising and bandwidth extension. We also proposed a shared SFI-STFT discriminator for unified adversarial training across diverse sampling rates. Together, these results establish TF-Restormer as a universal and practical framework for robust speech restoration across diverse sampling conditions.
Reproducibility statement
We make our implementation code and demo anonymously accessible. Our anonymous link for pre-trained TF-Restormer and inference code is also included in demo pages: https://tf-restormer.github.io/demo
Ethics statement
We use only public speech corpora and collect no new personal data. We do not attempt speaker re-identification, and we do not redistribute raw audio. Aware of potential misuse (e.g., covert monitoring), we will apply access controls and intended-use restrictions and require legal compliance for any release.
Appendix A Use of Large Language Models
This paper was written by the authors. Large language models (LLMs) were used only for minor language polishing.
Appendix B Details of Training Procedure
B.1 Simulation of training dataset
Clean speech source
The model is trained with VCTK training set. VCTK corpus (Yamagishi et al., 2019) is a multi-speaker English corpus containing 110 speakers with different accents. We split it into a training part VCTK-Train and a testing part VCTK-Test. The version of VCTK we used is 0.92. To follow the data preparation strategy of previous restoration studies Liu et al. (2022b), only the  mic1 microphone data is used for experiments, and p280 and p315 are omitted for the technical issues. For the remaining 108 speakers, the last 8 speakers, p360,p361,p362,p363,p364,p374,p376,s5 are split as test set VCTK-Test for super-resolution subtask. Within the other 100 speakers, p232 and p257 are also excluded because they are used in the test set VCTK-ND and VCTK+DEMAND datasets. Therefore, the remaining 98 speakers are used as training data.
Simulation pipeline
To simulate input signal for training, we randomly applied the various distortions based on the pipeline as shown in Figure 4. In particular, we sequentially applied physical and digital distortions. The physical distortions include convolution of transfer function mainly caused by reverberation at indoor environment. We used RIR samples from DNS dataset (Reddy et al., 2022). Note that we compensated time-delay effect from the convolution by applying direct component of RIR to the corresponding target speech signal. Then, as a second physical distortion, we added various background and interfering noises using noise samples (Reddy et al., 2022) and simulated colored gaussian noise. Each noise source is independently applied with signal-to-noise (SNR) ratio ranging from 0 to 20 dB. Then, as a final stage of physical distortion, we applied band pass filtering (BPF) to account for the recording condition of microphone such as occlusion, hardware properties, in this study, we mainly considered occlusion effect for the simulation. Also, to remove the phase distortion from the BPF, we applied as zero-phase filtering because the model does not need to consider these effect, only to make the learning process complicated. As a final step for physical simulation, we randomly scaled the level of signals from -35 to -15 dB Full Scale (dBFS). We also scaled the speech sources along with the corresponding input.
Figure 4: Noisy-distorted speech input simulation pipeline. The simulation procedure is partitioned to physical distortion and digital distortion.
Then, three kinds of digital distortions were simulated in sequence. We randomly applied audio clipping, crystalizer, flanger, and crusher as digital effects, each introducing characteristic nonlinear saturation, spectral over-enhancement, comb filtering, or quantization noise (detailed parameter ranges are summarized in Table 5). Afterward, digital codec compression was applied to emulate transmission artifacts, using either MP3 or OGG (Vorbis/Opus) encoding. Finally, the processed signals were randomly downsampled to 8 or 16 kHz to simulate low-bandwidth recording and communication scenarios.
Table 5: List of applied distortions with probabilities and parameter ranges.
Augmentation Prob. Param. name Range / Values Notes
RIR convolution 0.50 - - direct-path delay compensated
Sample Noise 1.00 SNR (dB) $[0,20]$ from DNS dataset
Colored Gaussian Noise 1.00 SNR (dB) $[0,20]$
exponent $\beta$ $[0.75,1.5]$
Band-limiting (BPF) 0.5 $f_{1}$ (Hz) $[500,1500]$ zero-phase
(occlusion FIR) $f_{2}$ (Hz) $f_{1}+[200,500]$ transition band upper edge
cut_gain $(0.1,0.3)$ stopband gain, applied as $g^{\beta}$
$\beta$ $[0.25,1.00]$ (thus effective stopband $\approx[0.22,0.55]$)
taps odd in $[31,61]$ firwin2, $f_{s}{=}16$k ( $f_{N}{=}8$k )
Clipping 0.5 level (dB) [15,0][-15,0] hard clipping threshold
Crystalizer 0.15 intensity $[1,4]$ spectral “sharpening”
Flanger 0.05 depth $[1,5]$ short-delay comb filtering
Crusher (bit-depth) 0.10 bits $[1,9]$ quantization/aliasing
Codec (any) 0.30 one of the following
   MP3 bit rate (kbps) $[4,16]$ variable bit-rate sampled uniformly
   OGG encoder vorbis, opus random choice
Frequency Masking 1.00 $F_{bw}$ (freq. bins) $[0,10]$
# masks $[0,3]$ set to $[0,1]$ in adversarial training
Time Masking 1.00 $T_{dur}$ (frames) $[0,10]$
# masks $[0,2]$ set to $[0,1]$ in adversarial training
Downsample 1.00 target $f_{s}$ {8k (0.25), 16k (0.75)}
B.2 Training Details
For pretraining, TF-Restormer was optimized with batch size set to 2 to train on a single GPU of NVIDIA RTX 6000 Ada Generation 48GB and AdamW (Loshchilov & Hutter, 2019) as our optimizer. For pretraining, we update 200,000 steps with VCTK dataset. The length of training utterance is set to 3 seconds. Then, adversarial training is followed with 200,000 steps. For both the stages, we use learning rate 2.0e-4 with betas (0.9, 0.995) and learning rate decay 0.9 at every 10,000 steps after 100,000 steps for pretraining whereas 10,000 steps for adversarial training.
Each stage is trained with a linear warm-up for 5,000 steps for the generator. On the other hand, the discriminators in the adversarial training perform 2 optimizations for every single generator’s optimization step without warm-up. For the discriminators, the AdamW optimizer is used with betas (0.8, 0.999). Based on conventional multi-scale STFT discriminators (Défossez et al., 2023), we used the proposed multi-scale SFI-STFT discriminator with STFT window durations of [20, 40, 60, 80, 100] ms, which allows for the capture of spectral information at different resolutions.
Appendix C Details of Model Configuration
For TF-Restormer, $C_{E}$ and $B_{E}$ for encoder are set to 128 and 6 while $C_{D}$ and $B_{D}$ are set to 64 and 3. The kernel size in ConvFFN and the number of heads in MHSA/MHCA are commonly set to $K=7$ and $H=4$, respectively. For frequency projection layer, $F_{\text{proj}}$ is set to 512.
For offline TF-Restormer, each input mixture is normalized by dividing it by its standard deviation and the enhanced output is rescaled by the same factor. For streaming version of TF-Restormer, two mamba blocks are used in the time module with $d_{\text{state}}=16$, causal Conv1D kernel size $3$ with expansion factor $4$. For streaming version, we still use the non-causal Conv2D layer for input and output projection for robust restoration, therefore the latency increases by two frames, total latency of 80ms (40 ms window, 20 ms hop). Overall, the model size of TF-Restormer is 30.1M for offline mode and 19.0M for streaming mode, which are smaller sizes compared to the existing models.
C.1 Comparison of the model size and RTF
In Table 6, we compare model size and multiply-accumulate operations (MACs) for a 1-second-long input using ptflops package [^2] https://github.com/sovrasov/flops-counter.pytorch. We also measure real-time factor (RTF) measured on 4-second-long samples with an NVIDIA RTX 4090. Conventional models operate at fixed input-output sampling rates, which results in fixed MACs regardless of the task configuration. In contrast, TF-Restormer adapts its computation depending on the input and output rates $f_{E}$ and $f_{D}$.
Among baselines, StoRM requires 50 diffusion steps, leading to very high MACs and RTF despite its moderate model size. UNIVERSE and UNIVERSE++ reduce the number of steps (8 by default in the open implementation), which lowers the runtime cost compared to StoRM, but their model sizes remain relatively large and the diffusion process cannot be adapted for streaming, representing a fundamental limitation. TF-Locoformer, built on a dual-path design, involves higher computational complexity but benefits from effective parallelism, so its RTF is not as large as its MACs might suggest; its parameter size is also smaller than most diffusion- or vocoder-based systems.
Our proposed TF-Restormer also follows a dual-path formulation, so the raw MACs are relatively large. Nevertheless, RTF remains low in practice, comparable to or even faster than prior dual-path models. Crucially, TF-Restormer optimizes computation according to the input and output sampling rates: for instance, in the $8\rightarrow 16$ kHz setting, redundant high-frequency processing is skipped, yielding a very low RTF. Also, the streaming variant maintains consistently low RTF while preserving accuracy, demonstrating its suitability for real-time applications.
[…]
Model Model Size (M) $f_{E}\to f_{D}$ (kHz) MACs(G) RTF
VoiceFixer 70.3 44.1 $\to$ 44.1 12.9 0.010
StoRM 55.1 16 $\to$ 16 156.4 0.520
[…]
TF-Locoformer 14.9 16 $\to$ 16 246.9 0.025
48 $\to$ 48 731.6 0.088
TF-Restormer 30.1   8 $\to$ 16 240.8 0.009
  8 $\to$ 44.1 308.4 0.017
16 $\to$ 16 440.9 0.034
16 $\to$ 48 518.7 0.053
TF-Restormer-streaming 19.0   8 $\to$ 16 114.7 0.012
  8 $\to$ 44.1 138.1 0.018
16 $\to$ 16 214.5 0.035
16 $\to$ 48 242.0 0.049
Appendix D Simulation of VCTK Noisy Distorted Input
The noisy-distorted input from VCTK testset in Table 3 was generated by corrupting clean VCTK utterances with additive noise from DEMAND (Thiemann et al., 2013) and colored Gaussian noise, RIR samples from RWCP (Nakamura et al., 2000) and AIR (Jeub et al., 2009) for reverberation, and distortions such as clipping and band-limiting. Additional digital effects including audio codecs (MP3, OGG) were applied before resampling to various rates (8-48 kHz). This simulation aligns the training pipeline while maintaining samples partitioning of speech, noise, and RIR sources.
The details of parameter range are summarized in Table 7.
Table 7: List of applied distortions with probabilities and parameter ranges.
Augmentation Prob. Param. name Range / Values Notes
RIR convolution 0.50 - - direct-path delay compensated
Sample Noise 1.00 SNR (dB) $[5,20]$ from DNS dataset
Colored Gaussian Noise 1.00 SNR (dB) $[5,20]$
exponent $\beta$ $[0.75,1.5]$
Band-limiting (BPF) 0.20 $f_{1}$ (Hz) $[2000,4000]$ zero-phase
(occlusion FIR) $f_{2}$ (Hz) $f_{1}+[200,500]$ transition band upper edge
cut_gain $(0.1,0.3)$ stopband gain, applied as $g^{\beta}$
$\beta$ $[0.25,0.75]$ (thus effective stopband $\approx[0.22,0.55]$)
taps odd in $[31,61]$ firwin2, $f_{s}{=}16$k ( $f_{N}{=}8$k )
Clipping 0.20 level (dB) [10,0][-10,0] hard clipping threshold
Crystalizer 0.10 intensity $[1,2]$ spectral “sharpening”
Flanger 0.05 depth $[1,3]$ short-delay comb filtering
Crusher (bit-depth) 0.10 bits $[1,5]$ quantization/aliasing
Codec (any) 0.25 one of the following
   MP3 bit rate (kbps) $[16,64]$ variable bit-rate sampled uniformly
   OGG encoder vorbis, opus random choice
Frequency Masking 1.00 $F_{bw}$ (freq. bins) $[0,5]$
# masks $[0,1]$
Time Masking 1.00 $T_{dur}$ (frames) $[0,5]$
# masks $[0,1]$
Appendix E Ablation Details
E.1 Effects of scaled log-spectral loss
Figure 5: Gradient profiles of the proposed scaled log-spectral loss $\partial\ell/\partial d=w/(d+w)$ for different scale factors $w$. The curves show that the gradient is 11 near zero error and monotonically decreases as the distance d=|ys|d=|y-s| grows. Smaller $w$ values make the loss more sensitive to fine spectral deviations, while larger $w$ values maintain stronger gradients over broader error ranges.
Figure 5 illustrates the gradient behavior of the proposed scaled log-spectral loss, defined as $\partial\ell/\partial d=w/(d+w)$ where d=|ys|d=|y-s| denotes the spectral distance and $w$ is a scale factor. Unlike conventional $\ell_{1}$ or $\ell_{2}$ criteria, whose gradients are either constant regardless of error magnitude ($\ell_{1}$) or increase proportionally with larger errors ($\ell_{2}$), the proposed formulation yields gradients that are strongest near $d\approx 0$ and gradually diminish once $d$ exceeds $w$. This mechanism emphasizes regions where the spectrum is already well-aligned, thereby preserving fine details, while suppressing unstable updates from heavily corrupted regions. The figure shows that when $d=w$, the gradient magnitude stabilizes at $0.5$, providing a natural balance between emphasizing accurate components and de-emphasizing severely mismatched ones. As $w$ decreases, the loss becomes more sensitive to smaller deviations, further suggesting subtle spectral structures that are otherwise neglected in conventional losses.
E.2 Illustration of encoder-decoder structure ablation
Figure 6 provides detailed comparisons of the three decoder designs considered in our ablation.
(a) Decoder-only. This variant directly inserts extension queries into the decoder without an encoder counterpart. The decoder therefore bears the full burden of modeling both the observed input band and the missing high-frequency bands, resulting in heavier computation and weaker inductive bias from the input.
(b) Encoder-decoder without MHCA. Here the encoder first analyzes the input bandwidth, and the decoder has the same internal structure as the encoder but receives projected extension queries. Although this design separates analysis and reconstruction, the decoder relies only on self-attention within the extended sequence, and does not explicitly exploit encoder features for reconstruction.
(c) Encoder-decoder with MHCA (proposed). In our final design, the decoder additionally uses a frequency cross-self module, where encoder outputs serve as key-value inputs for cross-attention while extension queries act as queries. This enables direct conditioning of high-frequency synthesis on encoder features, while the self-attention within the decoder refines spectral structure among extended bins. As a result, the encoder specializes in processing the observed input, and the lightweight decoder focuses on plausible high-frequency generation guided by encoder information.
These illustrations highlight how the proposed encoder-decoder with MHCA achieves a clear division of labor: the encoder concentrates on input-bandwidth analysis, and the decoder selectively extends spectral content with cross-conditioning, leading to better efficiency and stability compared to the other two designs.
Figure 6: Unit modules in TF-Encoder and TF-Decoder. The (a) time module is based on MHSA with RoPE while (b) the frequency encoder module is based on MHSA with frequency projection layer. (c) The frequency decoder module utilize MHCA based on key/value from the encoder features
Appendix F Evaluation on Real-recorded Dataset
VoxCeleb Data
We test TF-Restormer on 50 selected utterances from VoxCeleb1 (Nagrani et al., 2017), balanced across gender and covering the Speech Transmission Index (STI) range 0.75-0.99, to examine performance on real-recorded speech with diverse speakers (Su et al., 2020). For VoxCeleb Data, our TF-Restormer shows as natural restoration outputs as recent competing vocoder or diffusion-based methods in terms of UTMOS, WVMOS, and DNSMOS.
Table 8: Comparison with prior work on Voxceleb real data.
VoxCeleb (HiFi-GAN-2 validation set, real data)
Model UTMOS ($\uparrow$) WVMOS ($\uparrow$) DNSMOS ($\uparrow$)
Input 2.76 $\pm$ 0.13 2.90 $\pm$ 0.16 2.72 $\pm$ 0.11
HiFi++ (Andreev et al., 2022) 2.76 $\pm$ 0.13 2.68 $\pm$ 0.14 2.98 $\pm$ 0.07
VoiceFixer 2.60 $\pm$ 0.09 2.79 $\pm$ 0.09 3.08 $\pm$ 0.06
DEMUCS (Défossez et al., 2019) 3.51 $\pm$ 0.08 3.72 $\pm$ 0.08 3.27 $\pm$ 0.04
StoRM 3.29 $\pm$ 0.08 3.54 $\pm$ 0.09 3.17 $\pm$ 0.04
BBED (Lay et al., 2023) 3.30 $\pm$ 0.10 3.47 $\pm$ 0.08 3.23 $\pm$ 0.04
HiFi-GAN-2 (Su et al., 2021) 3.67 $\pm$ 0.09 3.96 $\pm$} 0.06 3.32 $\pm$} 0.03
[…]
TF-Restormer 3.98 $\pm$ 0.09 3.82 $\pm$ 0.09 3.34 $\pm$ 0.06
TF-Restormer-streaming 3.64 $\pm$ 0.10 3.59 $\pm$ 0.12 3.29 $\pm$ 0.04
Appendix G Limitations and Future Works
Although TF-Restormer shows balanced improvements in signal fidelity and perceptual quality, several limitations remain. First, when the input speech is severely degraded, the high uncertainty in the observations can induce content or speaker artefacts. This reflects a common trade-off in speech restoration: adversarial training encourages perceptually natural reconstructions by sampling plausible high-frequency details, but can deviate from the exact clean reference, whereas purely supervised objectives stay closer to the ground truth yet tend to average out details, producing less natural-sounding results. Second, the model may exhibit language dependency. Because it is trained on extremely distorted scenarios with limited single-language datasets, the model (and WavLM used for perceptual loss) may amplify language-specific biases (e.g. memorizing certain phonetic patterns).
Finally, while the model targets universal speech restoration, it does not address multi-speaker scenarios such as separation or extraction. Future work includes mitigating artefacts under extreme degradations, extending training to multilingual and more diverse corpora, and adapting the framework to handle overlapped multi-speaker speech. Note that the TF dual-path architecture was originally proposed in speech separation, suggesting its potential extension to separation tasks.